# Definability of mad families of vector spaces and two local Ramsey theories

### v4

@@ 5.1. Block-splitting families
Now suppose that there is no $C\leq B$ which is almost disjoint from every element of $\mathcal{A}$ (if such a $C$ exists, we may apply the argument in the previous paragraph to $C$, and use the fact that $\mathcal{I}^{+}(\mathcal{A})$ is $\subseteq$-upward closed). By Lemma 4.11, let $\{A_{n}:n<\omega\}\subseteq\mathcal{A}$ be an infinite set and let $\{B_{n}:n<\omega\}$ be an infinite disjoint family such that $\langle B_{n}\rangle\subseteq\langle B\rangle\cap\langle A_{n}\rangle$ for all $n$. Let $\Gamma:\omega\to\omega$ be an infinite-to-one surjection. We define an infinite block sequence $(x_{n})_{n<\omega}$ inductively as follows: If $(x_{k})_{k<n}$ has been defined, let $x_{n}\in\langle B_{\Gamma(n)}\rangle$ be any vector $(x_{k})_{k<n}<x_{n}$. Having defined $(x_{n})_{n<\omega}$, we now define an interval partition $\mathcal{P}:=\{P_{m}:m<\omega\}$ as follows: Let $P_{0}:=\{0,\dots,\max(\operatorname{supp}(x_{n_{0}}))\}$, where $n_{0}$ is the least integer such that $\Gamma(n_{0})=0$. If $P_{m}$ has been defined with $\max(P_{m})=n_{m}$, Let $\min(P_{m+1})=n_{m}+1$, and let $\max(P_{m+1})$ be large enough so that for all $i\leq m+1$, there exists some $n$ such that $\Gamma(n)=i$ and $\operatorname{supp}(x_{n})\subseteq P_{m+1}$. This is possible as $\Gamma$ is infinite-to-one.
@@
Now let $\alpha<\mathfrak{bs}$ be such that $S_{\alpha}$ splits $\mathcal{P}$. For each $i$, let $T_{i}^{0}:=\{n<\omega:\Gamma(n)=i\wedge\operatorname{supp}(x_{n})\subseteq S_{\alpha}\}$ and $T_{i}^{1}:=\{n<\omega:\Gamma(n)=i\wedge\operatorname{supp}(x_{n})\cap S_{\alpha}=\emptyset\}$. By our construction of $\mathcal{P}$, we have that $T_{i}^{0}$ and $T_{i}^{1}$ are infinite for all $i$. By a similar reasoning to the first case, we may conclude that $\langle B\rangle\cap Z_{\alpha}$ and $\langle B\rangle\cap Z_{\alpha}^{c}$ are compatible with $A_{i}$ for all $i$, so they are both in $\mathcal{I}^{+}(\mathcal{A})$. ∎
@@
Let $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ be a family of subsets of vectors satisfying the conclusions of Lemma 5.5. We write $Z_{\alpha}^{0}:=Z_{\alpha}$ and $Z_{\alpha}^{1}:=Z_{\alpha}^{c}$ for all $\alpha$. With this, given any almost disjoint family $\mathcal{A}\subseteq E^{[\infty]}$ and $B\in\mathcal{H}^{-}(\mathcal{A})$, there exist $\alpha<\mathfrak{bs}$ and $\tau_{B}^{\mathcal{A}}\in 2^{\alpha}$ such that:
@@
Observe that for a fixed almost disjoint family $\mathcal{A}$, $\tau_{B}^{\mathcal{A}}\in 2^{<\mathfrak{bs}}$ is unique, and if $B^{\prime}\leq B$ and $B^{\prime}\in\mathcal{H}^{-}(\mathcal{A})$, then $\tau_{B}^{\mathcal{A}}\subseteq\tau_{B^{\prime}}^{\mathcal{A}}$. Let $\operatorname{Lim}(\mathfrak{c})$ be the set of limit ordinals below $\mathfrak{c}$. Let $\{Y_{\alpha}:\alpha<\operatorname{Lim}(\mathfrak{c})\}$ be an enumeration of $\mathcal{P}(E)$, and let $\{B_{\alpha}:\alpha<\operatorname{Lim}(\mathfrak{c})\}$ be an enumeration of $E^{[\infty]}$ in a way that for any $B\in E^{[\infty]}$ and $Y\subseteq E$, there are cofinally many $\alpha<\operatorname{Lim}(\mathfrak{c})$ such that $(Y_{\alpha},B_{\alpha})=(Y,B)$. We shall define our mad family $\mathcal{A}=\{A_{\alpha}:\alpha<\mathfrak{c}\}$ recursively, where $\mathcal{A}_{\alpha}:=\{A_{\beta}:\beta<\alpha\}$, along with a collection $\{A_{\alpha}^{\prime}:\alpha<\operatorname{Lim}(\mathfrak{c})\}$ of block sequences, and an injective sequence $\{\sigma_{\alpha}:\alpha<\operatorname{Lim}(\mathfrak{c})\}$ of elements of $2^{<\mathfrak{bs}}$ such that the following holds:
@@
    If $\beta<\alpha$ is a limit ordinal, then $\sigma_{\alpha}\not\sqsubseteq\sigma_{\beta}$.
@@
For our base case, we let $\mathcal{A}_{\omega}$ be any countable almost disjoint family. Assume that $\alpha\geq\omega$ is a limit ordinal, and that $A_{\beta}$, $A_{\beta}^{\prime}$ and $\sigma_{\beta}$ have been defined for $\beta<\alpha$, and satisfy the requirements stated above so far. We define the element $B:=B_{\alpha}$ depending on the following scenarios:
@@
    Suppose that $B_{\alpha}\in\mathcal{H}(\mathcal{A}_{\alpha})$, and:
@@
    Suppose that $B_{\alpha}\in\mathcal{H}(\mathcal{A}_{\alpha})$, and for all $B\leq B_{\alpha}$ which is almost disjoint from every element of $\mathcal{C}_{\alpha}$, there is no $B^{\prime}\leq B$ such that $\langle B^{\prime}\rangle\not\subseteq Y_{\alpha}$. Then we let $B:=B_{\alpha}$.
@@
    $C_{s^{\frown}0},C_{s^{\frown}1}\in\mathcal{H}^{-}(\mathcal{A})$ are such that $\langle C_{s^{\frown}0}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{0}$ and $\langle C_{s^{\frown}1}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{1}$ (which both exist by properties (ii) and (iii) of Lemma 5.5).

### v5

@@ 5.1. Block-splitting families
Now suppose that there is no $C\leq B$ which is almost disjoint from every element of $\mathcal{A}$ (if such a $C$ exists, we may apply the argument in the previous paragraph to $C$, and use the fact that $\mathcal{I}^{+}(\mathcal{A})$ is $\subseteq$-upward closed). By Lemma 4.13, let $\{V_{n}:n<\omega\}\subseteq\mathcal{A}$ be an infinite set and let $\{B_{n}:n<\omega\}$ be an infinite disjoint family such that $\langle B_{n}\rangle\subseteq\langle B\rangle\cap V_{n}$ for all $n$. Let $\Gamma:\omega\to\omega$ be an infinite-to-one surjection. We define an infinite block sequence $(x_{n})_{n<\omega}$ inductively as follows: If $(x_{k})_{k<n}$ has been defined, let $x_{n}\in\langle B_{\Gamma(n)}\rangle$ be any vector $x_{n-1}<x_{n}$. Having defined $(x_{n})_{n<\omega}$, we now define an interval partition $\mathcal{P}:=\{P_{m}:m<\omega\}$ as follows: Let $P_{0}:=\{0,\dots,\max(\operatorname{supp}(x_{n_{0}}))\}$, where $n_{0}$ is the least integer such that $\Gamma(n_{0})=0$. If $P_{m}$ has been defined with $\max(P_{m})=n_{m}$, Let $\min(P_{m+1})=n_{m}+1$, and let $\max(P_{m+1})$ be large enough so that for all $i\leq m+1$, there exists some $n$ such that $\Gamma(n)=i$ and $\operatorname{supp}(x_{n})\subseteq P_{m+1}$. This is possible as $\Gamma$ is infinite-to-one.
@@
Now let $\alpha<\mathfrak{bs}$ be such that $S_{\alpha}$ splits $\mathcal{P}$. For each $i$, let $T_{i}^{0}:=\{n<\omega:\Gamma(n)=i\wedge\operatorname{supp}(x_{n})\subseteq S_{\alpha}\}$ and $T_{i}^{1}:=\{n<\omega:\Gamma(n)=i\wedge\operatorname{supp}(x_{n})\cap S_{\alpha}=\emptyset\}$. By our construction of $\mathcal{P}$, we have that $T_{i}^{0}$ and $T_{i}^{1}$ are infinite for all $i$. By a similar reasoning to the first case, we may conclude that $\langle B\rangle\cap Z_{\alpha}$ and $\langle B\rangle\cap Z_{\alpha}^{c}$ are compatible with $V_{i}$ for all $i$, so they are both in $\mathcal{I}^{+}(\mathcal{A})$. ∎
@@
Let $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ be a family of subsets of vectors satisfying the conclusions of Lemma 5.4. We write $Z_{\alpha}^{0}:=Z_{\alpha}$ and $Z_{\alpha}^{1}:=Z_{\alpha}^{c}$ for all $\alpha$. With this, for a fixed almost disjoint family $\mathcal{A}\subseteq E^{[\infty]}$ and $B\in\mathcal{H}^{-}(\mathcal{A})$, there exist $\alpha<\mathfrak{bs}$ and $\tau_{B}^{\mathcal{A}}\in 2^{\alpha}$ such that:
@@
Observe that for a fixed almost disjoint family $\mathcal{A}$, $\tau_{B}^{\mathcal{A}}\in 2^{<\mathfrak{bs}}$ is unique, and if $B^{\prime}\leq B$ and $B^{\prime}\in\mathcal{H}^{-}(\mathcal{A})$, then $\tau_{B}^{\mathcal{A}}\subseteq\tau_{B^{\prime}}^{\mathcal{A}}$. Let $\operatorname{Lim}(\mathfrak{c})$ be the set of limit ordinals below $\mathfrak{c}$. Let $\{Y_{\alpha}:\alpha\in\operatorname{Lim}(\mathfrak{c})\}$ be an enumeration of $\mathcal{P}(E)$, and let $\{B_{\alpha}:\alpha\in\operatorname{Lim}(\mathfrak{c})\}$ be an enumeration of $E^{[\infty]}$ in a way that for any $B\in E^{[\infty]}$ and $Y\subseteq E$, there are cofinally many $\alpha\in\operatorname{Lim}(\mathfrak{c})$ such that $(Y_{\alpha},B_{\alpha})=(Y,B)$. We shall define our mad family $\mathcal{A}=\{A_{\alpha}:\alpha<\mathfrak{c}\}$ recursively, where $\mathcal{A}_{\alpha}:=\{A_{\beta}:\beta<\alpha\}$, along with a collection $\{A_{\alpha}^{\prime}:\alpha\in\operatorname{Lim}(\mathfrak{c})\}$ of block sequences, and an injective sequence $\{\sigma_{\alpha}:\alpha\in\operatorname{Lim}(\mathfrak{c})\}$ of elements of $2^{<\mathfrak{bs}}$ such that the following holds:
@@
    If $\beta<\alpha$ are limit ordinals, then $\sigma_{\alpha}\not\sqsubseteq\sigma_{\beta}$.
@@
For our base case, we let $\mathcal{A}_{\omega}$ be any countable almost disjoint family. Assume that $\alpha\geq\omega$ is a limit ordinal, and that $A_{\beta+n}$, $A_{\beta}^{\prime}$ and $\sigma_{\beta}$ have been defined for limit $\beta<\alpha$ and $n<\omega$, and satisfy the requirements stated above so far. We define the element $B$ depending on the following scenarios:
@@
    Suppose that $B_{\alpha}\in\mathcal{H}(\mathcal{A}_{\alpha})$ (so $B_{\alpha}\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ by Lemma 4.10), and:
@@
    Suppose that $B_{\alpha}\in\mathcal{H}(\mathcal{A}_{\alpha})$ (so $B_{\alpha}\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$), and for all $B\leq B_{\alpha}$ which is almost disjoint from every element of $\mathcal{C}_{\alpha}$, there is no $B^{\prime}\leq B$ such that $\langle B^{\prime}\rangle\not\subseteq Y_{\alpha}$. Then we let $B:=B_{\alpha}$.
@@
    $C_{s^{\frown}0},C_{s^{\frown}1}\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ are such that $\langle C_{s^{\frown}0}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{0}$ and $\langle C_{s^{\frown}1}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{1}$ (which both exist by properties (ii) and (iii) of Lemma 5.4).
We then define $\eta_{f}:=\bigcup_{n<\omega}\eta_{f\mathord{\upharpoonright}n}$ for any $f\in 2^{\omega}$, and note that $\eta_{f}\in 2^{<\mathfrak{bs}}$ as $\mathfrak{bs}$ has uncountable cofinality (as $\mathfrak{b}$ is regular and $\mathfrak{s}$ has uncountable cofinality, see Corollary 21 of [8]). We note the following observations:
  1. (1)
    If $s\sqsubseteq t$, then $C_{s}\leq C_{t}$ and $\eta_{s}\sqsubseteq\eta_{t}$.
  2. (2)
    If $s$ and $t$ are incompatible, then so are $\eta_{s}$ and $\eta_{t}$.
  3. (3)
    Consequently, if $f\neq g$, then $\eta_{f}$ and $\eta_{g}$ are incompatible.
Let $\sigma_{\alpha}:=\eta_{f}$. Observe that for all $\xi<\operatorname{dom}(\eta_{f})$, if $n<\omega$ is such that $\xi\in\operatorname{dom}(\eta_{f\mathord{\upharpoonright}n})$, and if $N$ is such that $D/N\leq C_{f\mathord{\upharpoonright}n}$, then:
$\mathcal{B}_{0}$ $:=\{A_{\beta+n}:\sigma_{\beta}\sqsubseteq\sigma_{\alpha}\text{ and }n<\omega\},$
$\mathcal{B}$ $:=\mathcal{B}_{0}\cup\mathcal{B}_{1}.$
Note that $\mathcal{B}\subseteq\mathcal{A}_{\alpha}$, so it is an almost disjoint family.
Claim 8.
$|\mathcal{B}|<\mathfrak{bs}$.
Proof.
It suffices to show that $|\mathcal{B}_{0}|<\mathfrak{bs}$ and $|\mathcal{B}_{1}|<\mathfrak{bs}$. Since $\operatorname{dom}(\sigma_{\alpha})<\mathfrak{bs}$, there are $\mathfrak{bs}$ many initial segments of $\sigma_{\alpha}$. This implies that there are $\mathfrak{bs}$ many $\beta<\alpha$ such that $\sigma_{\beta}\sqsubseteq\sigma_{\alpha}$, as if $\beta,\gamma<\alpha$ and $\beta\neq\gamma$, then $\sigma_{\beta}\neq\sigma_{\gamma}$ by the induction hypothesis. Therefore, $|\mathcal{B}_{0}|<\mathfrak{bs}$. $\mathcal{B}_{1}$ is a union of $\operatorname{dom}(\sigma_{\alpha})<\mathfrak{bs}$ many finite sets, so $|\mathcal{B}_{1}|<\mathfrak{bs}$ as well. ∎
By Lemma 2.3, the set:
$\{\langle B^{\prime}\rangle\cap\langle D\rangle:B^{\prime}\in\mathcal{B}\text{ and }\langle B^{\prime}\rangle\cap\langle D\rangle\text{ is infinite-dimensional}\}$
is an almost disjoint family of block subspaces of $\langle D\rangle$. Note that $\langle D\rangle$ is not in the set, as otherwise this implies that $\langle D\rangle\subseteq A_{\beta}$ for some $\beta<\alpha$, contradicting that $\mathcal{D}\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$. By Lemma 5.3, $|\mathcal{B}|<\mathfrak{bs}\leq\mathfrak{a}_{\mathrm{vec},\mathbb{F}}$, so there exists some $A_{\alpha}^{\prime}\leq D$ such that $A_{\alpha}^{\prime}$ is almost disjoint from every element of $\mathcal{B}$.
Claim 9.
  1. (1)
    For all $\xi\in\operatorname{dom}(\sigma_{\alpha})$, $\langle A_{\alpha}^{\prime}\rangle\setminus Z_{\xi}^{\sigma_{\alpha}(\xi)}$ is small.
  2. (2)
Proof.
  1. (1)
    By Lemma 4.8, $\langle A_{\alpha}^{\prime}\rangle\cap\langle A\rangle$ is very small for each $A\in\mathcal{F}_{\xi}$, so $\langle A_{\alpha}^{\prime}\rangle\cap Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small.
  2. (2)
    Let $\beta+n<\alpha$, where $\beta$ is a limit ordinal. If $A_{\beta+n}\in\mathcal{B}_{0}$, then $A_{\alpha}$ is almost disjoint from $A_{\beta+n}$ by construction, so we are done. Now assume that $A_{\beta+n}\notin\mathcal{B}_{0}$. Let $\xi$ be the least ordinal such that $\sigma_{\beta}(\xi)\neq\sigma_{\alpha}(\xi)$. By induction hypothesis, we have that $W:=\langle A_{\beta}^{\prime}\rangle\setminus Z_{\xi}^{\sigma_{\beta}(\xi)}=\langle A_{\beta}^{\prime}\rangle\setminus Z_{\xi}^{1-\sigma_{\alpha}(\xi)}$ is small (and hence very small). Since $A_{\beta+n}\leq A_{\beta}^{\prime}$, we have that for some small set $W^{\prime}$:
We complete the induction by letting $\{A_{\alpha+n}:n<\omega\}$ be any infinite disjoint family below $A_{\alpha}^{\prime}$. We note that $\mathcal{A}$ is a mad family — otherwise, if $B$ is almost disjoint from every element of $\mathcal{A}$, then $B\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ for all $\alpha<\mathfrak{c}$. Let $\alpha$ be any limit ordinal such that $B_{\alpha}=B$, and we have that $A_{\alpha}\leq B$, a contradiction. It remains to show that $\mathcal{A}$ is full.
To see that $\mathcal{A}$ is full, let $Y\subseteq E$, and suppose that $Y$ is $\mathcal{H}(\mathcal{A})$-dense below some $B\in\mathcal{H}(\mathcal{A})$. Since $\mathcal{H}(\mathcal{A})=\bigcup_{\alpha<\mathfrak{c}}\mathcal{H}(\mathcal{A}_{\alpha})$, we may let $\alpha$ be large enough so that $(Y_{\alpha},B_{\alpha})=(Y,B)$ and $B\in\mathcal{H}(\mathcal{A}_{\alpha})$. This means that in the induction step, we have either case (b) or (c). Observe that (c) is not possible — since $A_{\alpha}^{\prime}\leq B_{\alpha}$ is almost disjoint from every element of $\mathcal{C}_{\alpha}$ as $\mathcal{C}_{\alpha}\subseteq\mathcal{A}_{\alpha}$, and $A_{\alpha}^{\prime}\in\mathcal{H}(\mathcal{A})$ as $A_{\alpha}^{\prime}$ is compatible with $A_{\alpha+n}$ for all $n$. However, there is no $B^{\prime}\leq A_{\alpha}^{\prime}$ for which $\langle B^{\prime}\rangle\subseteq Y_{\alpha}$, contradicting that $Y$ is $\mathcal{H}(\mathcal{A})$-dense below $B$. This means that case (b) must have occurred in this induction step. Since $B_{\alpha^{\prime}}$ is compatible with $A_{\alpha+n}$ for all $n$, $B_{\alpha^{\prime}}\in\mathcal{H}(\mathcal{A})$ and $\langle B_{\alpha}^{\prime}\rangle\subseteq Y_{\alpha}$, as desired. ∎
5.3. Completely separable mad family of subspaces
We dedicate this section to defining a notion of completely separable almost disjoint families of subspaces and proving Theorem 1.6.
Definition 5.5.
An almost disjoint family $\mathcal{A}$ of subspaces is completely separable if for any subspace $U\subseteq E$, if $U\in\mathcal{I}^{+}(\mathcal{A})$ then there exists some $V\in\mathcal{A}$ such that $V\subseteq U$.
Lemma 5.6.
A completely separable almost disjoint family of subspaces is maximal.
Proof.
Suppose that $\mathcal{A}$ is a completely separable almost disjoint family of subspaces that is not maximal. Let $U\subseteq E$ be a subspace almost disjoint from every $V\in\mathcal{A}$. For any $V_{0},\dots,V_{n-1}\in\mathcal{A}$, we have that:
$U\setminus\bigcup_{i<n}V_{i}=U\setminus\bigcup_{i<n}U\cap V_{i}.$
By Lemma 4.8, $U\cap V_{i}$ is very small for all $i<n$, so $\bigcup_{i<n}U\cap V_{i}$ is very small. Since $U$ is big, we have that $U\setminus\bigcup_{i<n}V_{i}$ is big. Since $V_{0},\dots,V_{n-1}$ are arbitrary, $U\in\mathcal{I}^{+}(\mathcal{A})$. By the complete separability of $\mathcal{A}$, this implies that $V\subseteq U$ for some $V\in\mathcal{A}$, contradicting that $U$ and $V$ are almost disjoint. ∎
One may consider the alternative definition where we remove the subspace requirement of $U$ in Definition 5.5. However, it turns out that no such almost disjoint families of subspaces exist.
Proposition 5.7.
Let $\mathcal{A}$ be an infinite almost disjoint family of subspaces. There exists some $Y\in\mathcal{I}^{++}(\mathcal{A})$ such that for all $A\in\mathcal{A}$, $\langle A\rangle\not\subseteq Y$.
Proof.
Let $\{V_{n}:n<\omega\}\subseteq\mathcal{A}$. For each $n$, let $B_{n},C_{n}$ be block subspaces of $V_{n}$ such that $\langle B_{n}\rangle\cap\langle C_{n}\rangle=\emptyset$. Let $\Gamma:\omega\to\omega$ be an infinite-to-one surjection. We define a block sequence $(x_{m})_{m<\omega}$ as follows: Let $x_{0}\in\langle C_{\Gamma(0)}\rangle$ be any vector. If $x_{m}$ has been defined, let $x_{m+1}\in\langle C_{\Gamma(m+1)}\rangle$ be any vector such that $x_{m}<x_{m+1}$. We let $D_{n}:=(x_{m})_{m\in\Gamma^{-1}(n)}$, which is an infinite block sequence as $\Gamma$ is infinite-to-one. We write $D_{n}=(y_{k}^{n})_{k<\omega}$ Note that the block subspaces $D_{n}$ have the properties:
  1. (1)
    For all $z\in\langle D_{n}\rangle$, $\operatorname{supp}(z)\subseteq\bigcup_{k<\omega}\operatorname{supp}(y_{k}^{n})$.
  2. (2)
    If $m\neq n$, then $\bigcup_{k<\omega}\operatorname{supp}(y_{k}^{n})\cap\bigcup_{j<\omega}\operatorname{supp}(y_{j}^{m})=\emptyset$.
We define:
$Y:=\bigcup_{n<\omega}\langle D_{n}\rangle.$
Clearly $Y\in\mathcal{I}^{++}(\mathcal{A})$, as $\langle D_{n}\rangle\subseteq V_{n}\cap Y$ for all $n$.
Claim 10.
If $B\in E^{[\infty]}$ and $\langle B\rangle\subseteq Y$, then $B\leq D_{n}$ for some $n$.
Proof.
We write $B=(z_{k})_{k<\omega}$, and suppose that $z_{0}\in\langle D_{n}\rangle$. We shall show that $z_{k}\in\langle D_{n}\rangle$ for all $k$, which implies that $\langle B\rangle\subseteq\langle D_{n}\rangle$. Suppose that $z_{k}\in\langle D_{m}\rangle$ for some $m\neq n$ and $k>0$. $z_{0}+z_{k}\notin\langle D_{n}\rangle$, as:
$\operatorname{supp}(z_{0}+z_{k})=\operatorname{supp}(z_{0})\cup\operatorname{supp}(z_{k})\not\subseteq\bigcup_{j<\omega}\operatorname{supp}(y_{j}^{n}).$
Similarly, $z_{0}+z_{k}\notin\langle D_{m}\rangle$. For each $i\neq m,n$, since:
$\operatorname{supp}(z_{0})\cap\bigcup_{j<\omega}\operatorname{supp}(y_{j}^{i})=\emptyset,$
we may conclude that $z_{0}+z_{k}\notin\langle D_{i}\rangle$ either. Therefore, $z_{0}+z_{k}\in\langle B\rangle\setminus Y$, a contradiction. ∎
We may now finish the proof of the lemma. Suppose that there exists some $V\in\mathcal{A}$ such that $V\subseteq Y$. By the claim above, we have that $V\subseteq\langle D_{n}\rangle\subseteq V_{n}$ for some $n$. Since $\mathcal{A}$ is an almost disjoint family, we have that $V=V_{n}$. But this implies that $\langle B_{n}\rangle\subseteq V_{n}\subseteq\langle D_{n}\rangle\subseteq\langle C_{n}\rangle$, which contradicts that $\langle B_{n}\rangle\cap\langle C_{n}\rangle=\emptyset$. ∎
Proof of Theorem 1.6.
Let $\{Z_{\alpha}:\alpha<\mathfrak{bs}\}$ be a family of subsets of vectors satisfying the conclusions of Lemma 5.4. We write $Z_{\alpha}^{0}:=Z_{\alpha}$ and $Z_{\alpha}^{1}:=Z_{\alpha}^{c}$ for all $\alpha$. For a fixed almost disjoint family $\mathcal{A}\subseteq E^{[\infty]}$ and $B\in\mathcal{H}^{-}(\mathcal{A})$, there exist $\alpha<\mathfrak{bs}$ and $\tau_{B}^{\mathcal{A}}\in 2^{\alpha}$ such that:
  1. (1)
  2. (2)
Observe that for a fixed almost disjoint family $\mathcal{A}$, $\tau_{B}^{\mathcal{A}}\in 2^{<\mathfrak{bs}}$ is unique, and if $B^{\prime}\leq B$ and $B^{\prime}\in\mathcal{H}^{-}(\mathcal{A})$, then $\tau_{B}^{\mathcal{A}}\subseteq\tau_{B^{\prime}}^{\mathcal{A}}$. Let $\{U_{\alpha}:\alpha<\mathfrak{c}\}$ be an enumeration of all subspaces of $E$. We may need that for any subspace $U\subseteq E$, there are cofinally many $\alpha<\mathfrak{c}$ such that $U_{\alpha}=U$. We shall define our mad family $\mathcal{A}=\{A_{\alpha}:\alpha<\mathfrak{c}\}$ recursively, where $\mathcal{A}_{\alpha}:=\{A_{\beta}:\beta<\alpha\}$, along with an injective sequence $\{\sigma_{\alpha}:\alpha<\mathfrak{c}\}$ of elements of $2^{<\mathfrak{bs}}$ such that the following holds:
  1. (1)
  2. (2)
  1. (1)
    If $U_{\alpha}\in\mathcal{I}^{+}(\mathcal{A}_{\alpha})$, let $C_{\emptyset}\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ such that $\langle C_{\emptyset}\rangle\subseteq U_{\alpha}$. This is possible by Lemma 4.14. If $U_{\alpha}\notin\mathcal{I}^{+}(\mathcal{A}_{\alpha})$, let $C_{\emptyset}\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ be any block subspace.
  2. (2)
  3. (3)
    $C_{s^{\frown}0},C_{s^{\frown}1}\in\mathcal{H}^{-}(\mathcal{A}_{\alpha})$ are such that $\langle C_{s^{\frown}0}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{0}$ and $\langle C_{s^{\frown}1}\rangle\subseteq\langle C_{s}\rangle\cap Z_{\alpha_{s}}^{1}$ (which both exist by properties (ii) and (iii) of Lemma 5.4).